OpenStudy (anonymous):

Solve by using elimination (we'll pretend like this is all one bracket and it's a system): {4x-2y=3 {2y-4x=3/2

OpenStudy (ℐℵk℣ⱱøƴȡ◆):

4 x-2 y = 6 y-4 x = 3/2

OpenStudy (inkyvoyd):

Add the two equations.

OpenStudy (haseeb96):

4x-2y=3 .....eq#1 2y -4x =3/2 ......eq#2 is the equation #2 is in this form 2x -4y =3/2 please check it from yur note book

OpenStudy (anonymous):

add the two equations -> 0 = 9/2

OpenStudy (haseeb96):

yes i got but this question does not exist ...